Assuming the company estimates bad debts at an amount equal to 2% of credit sales.

The following information relates to a company's accounts receivable: gross accounts receivable balance at the beginning of the year, $300,000; allowance for uncollectible accounts at the beginning of the year, $25,000 (credit balance); credit sales during the year, $1,500,000; accounts receivable written off during the year, $16,000; cash collections from customers, $1,450,000. Assuming the company estimates bad debts at an amount equal to 2% of credit sales.


 1. Calculate bad debt expenses for the year.

 2. Calculate the year-end balance in the allowance for uncollectible accounts. 
Answer #1

1) Bad debt expense = 1500000*2% = 30000

2) Ending balance = 25000+30000-16000 = 39000